# A Dataset and Benchmark for Copyright Protection from Text-to-Image Diffusion Models

### v1

@@ Abstract
Copyright is a legal right that grants creators the exclusive authority to reproduce, distribute, and profit from their creative works. However, the recent advancements in text-to-image generation techniques have posed significant challenges to copyright protection, as these methods have facilitated the learning of unauthorized content, artistic creations, and portraits, which are subsequently utilized to generate and disseminate uncontrolled content. Especially, the use of stable diffusion, an emerging model for text-to-image generation, poses an increased risk of unauthorized copyright infringement and distribution. Currently, there is a lack of systematic studies evaluating the potential correlation between content generated by stable diffusion and those under copyright protection. Conducting such studies faces several challenges, including i) the intrinsic ambiguity related to copyright infringement in text-to-image models, ii) the absence of a comprehensive large-scale dataset, and iii) the lack of standardized metrics for defining copyright infringement. This work provides the first large-scale standardized dataset and benchmark on copyright protection. Specifically, we propose a pipeline to coordinate CLIP, ChatGPT, and diffusion models to generate a dataset that contains anchor images, corresponding prompts, and images generated by text-to-image models, reflecting the potential abuses of copyright. Furthermore, we explore a suite of evaluation metrics to judge the effectiveness of copyright protection methods. The proposed dataset, benchmark library, and evaluation metrics will be open-sourced to facilitate future research and application. The website and dataset can be accessed website / dataset.
@@ 1 Introduction
Figure 1: The pivotal role of datasets in copyright protection. CPDM dataset encompasses copyrighted images, their corresponding prompts, and images suspected of infringement generated by Stable Diffusion (SD) models using these prompts. Users can assess the potential for infringement of a generated image by directly comparing it with the copyrighted image using CPDM metrics. Additionally, we offer a benchmark for unlearning methods, providing a standardized testing ground, along with reference metrics for fair comparisons.
@@
Text-to-image generative model have recently emerged as a significant topic in computer vision, demonstrating remarkable results in the area of generative modelings [14, 34]. These models bridge the gap between language and visual content by generating realistic images from textual descriptions. However, rapid advancements in text-to-image generation techniques have raised concerns about copyright protection, particularly unauthorized reproduction of content, artistic creations, and portraits [4]. A specific concern arises from the use of Stable Diffusion (SD), a state-of-the-art text-conditional latent diffusion model, which has sparked global discussions on copyright, privacy, and safety [42, 9]. Current research lacks systematic evaluations of the correlation between Stable Diffusion model outputs and potential copyright infringements. This gap stems partly from the inherent complexity in delineating copyright infringement criteria for text-to-image generative models, further compounded by the absence of extensive inference datasets and standardized benchmarks for copyright infringement assessment. In this study, we introduce the first comprehensive dataset tailored for copyright protection in this domain, along with novel metrics to evaluate the efficacy of various copyright protection strategies.
Figure 2: Examples of CPDM dataset composition and unlearned results for copyright protection. The dataset primarily comprises four categories: Style, Portrait, Artistic Creation Figure, and Licensed Illustration. The illustrations are exclusively provided and authorized by the Anonymous Artist, while the remaining images are sourced from WikiArt, WikiArtPedia, and the Internet. The unlearned images are obtained using the methods described in the benchmark Sec. 5.
@@
Initially, it is crucial to define what constitutes copyright infringement in content produced by text-to-image generative models [38]. In this study, we focus on infringement within 2D artistic works. Drawing on expertise from copyright protection specialists, including artists and lawyers, we contend that a unique painting style of an artist, virtual representations in artistic creations, and individual portraits all represent forms of creative expression deserving of legal protection. In order to identify instances of infringement in these contexts, a comprehensive analysis encompassing both technical and semantic aspects of the generated content is indispensable.
In light of the aforementioned considerations, we have developed evaluation metrics designed to identify potential copyright infringements in generated content and to assess the efficacy of unlearning methods aimed at making generative models (i.e. Stable Diffusion) forget what it has learned from the potentially infringed images [3, 12, 16, 15, 22]. To establish the validity of these metrics and enable equitable comparisons across different unlearning approaches, we introduce a novel dataset: Copyright Protection from Diffusion Model (CPDM) dataset. This dataset encompasses a collection of images with their original copyright attributes, associated prompts for text-to-image generation via Stable Diffusion, and outputs indicative of potential infringement. Additionally, we have classified these images into four distinct categories to ensure a comprehensive and instructive representation, as illustrated in Fig. 2.
[…]
Potential for Infringement
[…]
Effectiveness of Unlearning
[…]
Extent of Model Degradation during Unlearning
[…]
Moreover, we performed comprehensive benchmark tests on the newly proposed CPDM dataset. In our experiments, we utilized gradient ascent-based and response-based pruning methods for unlearning, as comparison baselines for other unlearning approaches, specifically targeting the Stable Diffusion model. Our evaluation methodology for these unlearning approaches centers on the accuracy of targeted concept removal pertaining to copyright attributes, examining this process from two principal perspectives:
@@
Effectiveness of Unlearning. We quantify the similarity between the original copyright images and their unlearned counterparts on feature level, utilizing our proposed CM metric. Additionally, we evaluate changes of CLIP scores for text-image similarity, indicating the extent to which the prompt triggering potential infringement is neutralized.
@@
Extent of Model Degradation during Unlearning. The unlearning process inherently degrades the model by eliminating certain infringement-suspected concepts. Nevertheless, it is vital to preserve the Stable Diffusion model’s ability for copyright-irrelevant generation. We assess the degree of model degradation using the widely-recognized FID (Fréchet Inception Distance) metric [20].
@@
For a detailed analysis, see Tab. 1. This evaluation provides valuable insights into assessing copyright infringement and the efficacy of unlearning methods in reducing infringement risks, while preserving the ability to generate non-infringing content. Additionally, our benchmark facilitates a direct comparison of various methods.
@@ 2.1 Text-to-image Generative Models
Recently text-to-image diffusion models [34] have emerged as a crucial research area attracting wide attention. These state-of-the-art methods [30, 34, 35, 32, 2] have exhibited remarkable capabilities in transforming textual information into visually coherent and realistic images, often demonstrating high performance in terms of accuracy. The advancements in these techniques have opened up a plethora of possibilities for a wide range of downstream tasks, such as image editing [24, 13, 6, 1], image denoising [21, 41] and super-resolution [25, 10], etc. The advancements in text-to-image generative models have significantly impacted various industries. However, they also pose challenges to copyright protection. As these models become increasingly adept at creating high-quality images, distinguishing between original artworks and generated ones becomes more complex. This convergence raises critical questions about authorship, intellectual property rights, and the implications for plagiarism in the digital era. Addressing these concerns is imperative.
2.2 Model Unlearning
@@
[4] highlights that the privacy of diffusion models is significantly lower compared to generative adversarial networks (GANs)  [14]. Under the diffusion framework, models tend to retain certain images from the training data, potentially generating outputs that closely resemble the original images. To remove explicit artwork from large models,  [9] presents a fine-tuning method for concept removal from diffusion models. Additionally,  [42] presents the “Forget-Me-Not” method, which enables the targeted removal of specific objects and content from large models within a span of 30 seconds, while minimizing the impact on other content.  [38] explores whether diffusion models create unique artworks or directly replicate certain content from the training dataset during image generation. Furthermore, there exist numerous model unlearning methods in the context of image-related tasks, as evidenced by  [3, 12, 16, 15, 22], among others. Regarding the comprehensive review of model unlearning, it has provided an overview of unlearning algorithms in various domains such as images, tables, text, sentences, and graphs  [43, 29]. Although machine unlearning is designed to protect the privacy of target samples, the paper  [5] has demonstrated, in the context of model classification tasks, that machine unlearning might leave traces. Consequently, it is imperative to consider these risks when developing unlearning algorithms for text-to-image generation models.
2.3 Metrics for Images Similarity
Accurately measuring image similarity has always been a challenge yet to be perfectly addressed. For image denoising and super-resolution tasks, researchers have introduced evaluation metrics such as Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index Measure (SSIM) to assess the fine-grained similarity between two images [40]. However, these metrics are limited in their ability to evaluate only near-identical images and lack the capacity to assess higher-level similarities. For specific generative tasks, Fréchet Inception Distance (FID)  [19] has become a prevalent metric, spanning from GAN models to diffusion methods. Nonetheless, FID is designed primarily to evaluate the distance between two sets of images, typically using realistic images as reference points. Consequently, it offers limited insight when dealing with two specific images and their inherent similarity.
2.4 Works in Artistic Image Communities
@@
With the emergence of painting capabilities in models like Stable Diffusion, there has been a growing surge in activity and attention within communities focused on image copyright protection. For instance, websites such as stablediffusion.fr/artists and urania.ai/top-sd-artists have gained prominence. These platforms have curated collections of images that are stylistically similar to the work of over a thousand artists, both contemporary and classical. These artists’ styles can be imitated using Stable Diffusion models. In comparison to the efforts of these image communities, our approach involves the collection of real, valuable, and specific images from the art world to be used as training examples for image generation models. This represents a more rigorous form of style imitation. In contrast, the artistic images in the provided links tend to focus on capturing certain aspects of an artist’s style, fitting into a broader category of style imitation. For details, see the supplementary materials on page 5.
2.5 Policy, Legal, and Social Impact
@@
The increasing global popularity of AIGC highlights the importance of privacy and copyright issues. AI companies, including OpenAI, have taken measures to address concerns related to data security. The US has proposed establishing a new government agency responsible for approving large-scale AI models. Furthermore, the Chinese Cyberspace Administration has published a document emphasizing AIGC security issues [^1] http://www.cac.gov.cn/2023-04/11/c_1682854275475410.htm. Recently enacted legislation, such as the General Data Protection Regulation (GDPR [^2] https://gdpr-info.eu/) in the European Union, the California Consumer Privacy Act (CCPA [^3] https://oag.ca.gov/privacy/ccpa) in California, and the Personal Information Protection and Electronic Documents Act (PIPEDA [^4] https://laws-lois.justice.gc.ca/ENG/ACTS/P-8.6/index.html) in Canada, have legally solidified this right [43, 5].
3 CPDM Metrics
Determining whether two works constitute plagiarism has long been a pressing issue in both the arts and legal domains. Perceptual evaluation metrics based on the human visual system, such as LPIPS [44], have achieved evaluation results that align more closely with human perception. By employing deep feature measurements to assess image similarity, these metrics produce perceptually accurate evaluations. Besides, in the realm of videos, perceptual evaluation metrics like VMAF [37] combine human visual modeling with machine learning techniques to achieve impressive results. However, previous research has been limited in addressing this challenge, particularly when measuring the similarity between anchor and generated images.
Both statements highlight the current research gap in copyright protection and the definition of similarity, underscoring the need for further investigation. Then, we collaborated with an artist, Anonymous Artist, who is currently active in the art industry and is notably distinguished in the field of illustration. This artist offered invaluable insights from a professional standpoint to help assess whether two images constitute plagiarism, considering elements such as brushstrokes, color palettes, lighting effects, and composition. For a thorough analysis, we divide these measurements into different categories: semantic and stylistic components. Our objective is to develop a formula that combines both components and provides a scalar metric to quantify the similarity between two images.
3.1 Semantic Metric
We leverage the CLIP[31] model to generate the semantic embedding, and calculate the metrics by:
[…]
$Loss_{sem}=MSE(emb_{ori}\,,\;emb_{gen})$ (2)
where $Image_{ori}$ and $Image_{gen}$ denote the anchor image and generated image respectively; $CLIP$ denotes the CLIP’s image encoder. In previous studies, cosine similarity has been predominantly employed as the evaluation metric, while, in this research, we utilize Mean Squared Error (MSE) instead. This decision is primarily motivated by two factors: first, the range of the MSE is significantly broader than that of cosine similarity, which makes it easier to observe changes resulting from unlearning; second, the adoption of MSE aligns better with the subsequent style metrics discussed below.
3.2 Style Metric
It is relatively more difficult to measure the similarity in style, in this part, inspired from the method in[11], we use the activation output by the CNN networks to calculate the features correlations given by the Gram matrix, in our work we leverage the InceptionV3[39], following the Fréchet Inception Distance (FID) metric:
[…]
$Loss_{style}=\sum_{i=1}^{n}w^{l}D^{l}$ (4)
where $Inception(Image,\,l)$ denotes passsing the $Image$ through an Inception network and extracting the feature maps from layer $l$. The Gram matrix is then computed to provide a style representation of the image at layer $l$. Furthermore, the dissimilarity between the original and generated images in each layer is represented by MSE of the Gram matrices in each corresponding layer. The total style metric, as described above, is determined by weighting factors $w^{l}$, which represent the contribution of each layer to the overall style metrics. In our work, $n$ is set to $4$, because there are four stages of the InceptionV3 model. The values of parameters $w^{l}$ needs to be fine-tuned according to the distribution of the images. We provide concrete values applicable to our dataset under Fig. 4. Finally, we denote the total metric as:
$CPDM$ $=(Loss_{sem}\times Loss_{style})^{2}$ (5)
Here, we adopt the squared term to emphasize the significant change before and after the unlearning process as well. And the effectiveness of this quantifiable metrics will be verified in Sec. 6.2.
Semantic metrics encapsulate the primary content information of images. Irrespective of the image’s stylistic attributes, its principal subject is extracted within the semantic metric. Furthermore, style metrics encompass attributes, including brushstroke depth, line thickness, color schemes, and compositional elements. These style metrics effectively represent a diverse array of image categories.
@@ 4 CPDM Dataset
4.1 Pipeline for Dataset Creation
Figure 3: Pipeline for the CPDM Dataset Creation. The CLIP Interrogator is utilized to convert copyrighted images into corresponding textual information. This text is subsequently refined and transformed into prompts, which are then inputted into a diffusion model to generate the corresponding infringing images.
We propose a pipeline to coordinate CLIP, ChatGPT, and diffusion models to generate a dataset that contains anchor images, corresponding prompts, and images generated by text-to-image models, reflecting the potential abuses of copyright. Initially, we collect a set of images that potentially contain copyrighted content, which serves as anchor images. Subsequently, these images are fed into the CLIP-interrogator, allowing us to obtain prompts that correspond to each anchor image. Finally, the prompts are used as input for the stable diffusion model, resulting in the generation of images by the stable diffusion model. Through manual comparisons, we assess whether there is evidence of copyright infringement in terms of style and semantics between the anchor images and the generated images. Ultimately, the anchor images, their corresponding prompts, and the images generated by the stable diffusion model constitute the core components of our dataset. This carefully curated dataset allows for a comprehensive examination of the copyright-related characteristics of the generated content, facilitating rigorous evaluations and analyses of the performance of various techniques in detecting and addressing copyright infringement.
4.2 Composition of the Dataset
Style Painting artworks often embody the distinctive style of the artist, encompassing aspects such as brushstrokes, lines, colors, and compositions. This artistic style is also a form of copyright that requires protection. WikiArt is an online user-editable visual art encyclopedia, a source from which numerous art-related datasets [23, 26] have been curated. WikiArt already features some 250,000 artworks by 3,000 artists, localized on 8 languages. These artworks are in museums, universities, town halls, and other civic buildings of more than 100 countries. We selected approximately 1500 artworks from 100 artists on WikiArt as the source of anchor images for prompt generation and corresponding content generation using stable diffusion.
Portrait The right of portrait refers to an individual’s control and use of their own portrait, including facial features, image, and posture. The purpose of publicity rights is to safeguard an individual’s privacy, personal dignity, and image integrity, preventing unauthorized use, disclosure, or alteration of their portrait. This legal protection aims to preserve an individual’s control over how their likeness is portrayed and commercialized. We utilized web scraping techniques to collect over 200 portrait images from Wikipedia, which is a free, web-based, multilingual encyclopedia that contains articles on a wide range of topics.
Artistic Creation Figure Artistic creations, including characters from animations and cartoons, are often protected by law. In this context, we refer to this category as ”artistic creation figures.” Similar to portraits, we have curated a dataset of 200 influential animated characters and figures by collecting information from reputable sources such as Wikipedia.
Licensed Illustration We have obtained authorization to use a portion of Anonymous Artist’s artworks in this study. Therefore, we can use his/her illustrations as part of the training dataset for fine-tuning stable diffusion, which will be utilized for the process of simulating infringing artistic paintings.

### v3

# A Dataset and Benchmark for Copyright Infringement Unlearning from Text-to-Image Diffusion Models

@@ Abstract
Copyright law confers upon creators the exclusive rights to reproduce, distribute, and monetize their creative works. However, recent progress in text-to-image generation has introduced formidable challenges to copyright enforcement. These technologies enable the unauthorized learning and replication of copyrighted content, artistic creations, and likenesses, leading to the proliferation of unregulated content. Notably, models like stable diffusion, which excel in text-to-image synthesis, heighten the risk of copyright infringement and unauthorized distribution.Machine unlearning, which seeks to eradicate the influence of specific data or concepts from machine learning models, emerges as a promising solution by eliminating the “copyright memories” ingrained in diffusion models. Yet, the absence of comprehensive large-scale datasets and standardized benchmarks for evaluating the efficacy of unlearning techniques in the copyright protection scenarios impedes the development of more effective unlearning methods. To address this gap, we introduce a novel pipeline that harmonizes CLIP, ChatGPT, and diffusion models to curate a dataset. This dataset encompasses anchor images, associated prompts, and images synthesized by text-to-image models. Additionally, we have developed a mixed metric based on semantic and style information, validated through both human and artist assessments, to gauge the effectiveness of unlearning approaches. Our dataset, benchmark library, and evaluation metrics will be made publicly available to foster future research and practical applications (website / dataset).
@@ 1 Introduction
Figure 1: The pipeline of Copyright Unlearning and the crucial role of Datasets and Benchmarks.
@@
Text-to-image generative models have recently emerged as a significant topic in computer vision, demonstrating remarkable results in the area of generative modeling [14, 35]. These models bridge the gap between language and visual contents by generating realistic images from textual descriptions. However, rapid advancements in text-to-image generation techniques have raised concerns about copyright protection, particularly unauthorized reproduction of content, artistic creations, and portraits [4]. A specific concern arises from the use of Stable Diffusion (SD), a state-of-the-art text-conditional latent diffusion model, which has sparked global discussions on copyright.
Machine Unlearning (MU), which aims to eliminate the influence of specific target data or concepts, presents a promising solution to the aforementioned challenges. Several studies have provided intuitive evidence of MU’s effectiveness in erasing the memory of copyrighted material from original models [42, 7, 26, 8]. However, the current body of research is limited by a lack of quantitative and systematic assessments that evaluate the extent to which MU can reduce the risk of copyright infringement. This limitation hinders the ability to make meaningful comparisons between existing MU approaches. The challenge is exacerbated by the inherent complexity of defining copyright infringement criteria for text-to-image generative models, as well as the scarcity of comprehensive inference datasets and standardized benchmarks for assessing copyright infringement. The absence of extensive copyright datasets obstructs researchers’ efforts to fully comprehend the copyright infringement risks associated with generative models. This, in turn, restricts their capacity to develop superior MU algorithms capable of addressing the legal risks effectively. As shown in Fig. 1, datasets and benchmarks are essential for forgetting copyrighted content and evaluating unlearning methods.
@@
Initially, it is crucial to define what constitutes copyright infringement in contents produced by text-to-image generative models [38]. In this study, we focus on infringement within 2D artistic works. Drawing on expertise from copyright protection specialists, including artists and lawyers, we contend that a unique painting style of an artist, virtual representations in artistic creations, and individual portraits all represent forms of creative expression deserving of legal protection. In order to identify instances of infringement in these contexts, a comprehensive analysis encompassing both technical and semantic aspects of the generated contents is indispensable.
In light of the aforementioned considerations, we introduce the first comprehensive dataset tailored for copyright protection in this domain: Copyright Protection from Diffusion Model (CPDM) dataset. Specifically, we have curated copyright images from four distinct categories that are most commonly suspected of infringement, as demonstrated in Fig. 2. This dataset encompasses a collection of original copyright images, associated prompts for text-to-image generation via Stable Diffusion, and indicatives for a series of features, represented by: 1) Potential for Infringement; 2) Effectiveness of Unlearning; 3) Extent of Model Degradation during Unlearning.
Potential for Infringement. This is quantified by feature-level similarities between the original copyright images and potentially infringing generated contents, denoted as CPDM metric (CM).
@@
Effectiveness of Unlearning. We reuse the CM metric to indicate the similarities between the original copyright images and their unlearned counterparts after processed by unlearning methods. Additionally, we evaluate changes of CLIP scores, denoted as $\Delta$CLIP, for text-image similarity. This indicates the extent to which the prompt that generates potential infringement is nullified.
@@
Extent of Model Degradation during Unlearning. The unlearning process inherently degrades the model by eliminating certain infringement-suspected concepts. Nevertheless, it is vital to preserve the Stable Diffusion model’s generation capacities for copyright-irrelevant contents. We assess the degree of model degradation using the widely-recognized FID (Fréchet Inception Distance) metric [20].
@@
Our benchmark facilitates a straightforward evaluation of potential copyright infringement, while facilitating comparison among various unlearning methods. Moreover, we perform comprehensive benchmark tests on our proposed CPDM dataset. In our experiments, we utilize gradient ascent-based and response-based pruning methods for unlearning, as comparison baselines for other unlearning approaches, specifically targeting the Stable Diffusion models. This evaluation provides valuable insights into assessing copyright infringement and the efficacy of unlearning methods in reducing infringement risks, while preserving the ability to generate non-infringing contents.
Figure 2: Examples of CPDM dataset composition and unlearned results for copyright protection.
@@ 2 Background and Related Work
Text-to-image Generative Model. Recently text-to-image diffusion models [35] have emerged as a crucial research area attracting wide attention. These state-of-the-art methods [31, 35, 36, 33, 2] have exhibited remarkable capabilities in transforming textual information into visually coherent and realistic images, often demonstrating high performance in terms of accuracy. The advancements in these techniques have opened up a plethora of possibilities for a wide range of downstream tasks, such as image editing [25, 12, 6, 1], image denoising [21, 41] and super-resolution [28, 9], etc. The advancements in text-to-image generative models have significantly impacted various industries. This convergence raises critical questions about authorship, intellectual property rights, and the implications for plagiarism in the digital era. Addressing these concerns is imperative.
Image Similarity Measurement. Determining plagiarism, a crucial concern in arts and legal domains, often relies on assessing image similarity. Existing metrics like PSNR and SSIM are limited to near-identical images, lacking the capacity to evaluate higher-level similarities [40]. Perceptual metrics like LPIPS, while aligning closely with human perception, may have limitations in capturing certain nuances[44]. In the realm of videos, metrics like VMAF, despite combining human visual modeling with machine learning techniques, may encounter challenges in specific scenarios[37]. While FID is prevalent in generative tasks, it primarily evaluates the distance between sets of images, leaving a gap in measuring inherent similarity between anchor and generated images [19]. This underscores the need for further investigation into copyright protection and similarity definition.
@@
Unlearning Method. Carlini et al.  [4] highlights that the privacy of Stable Diffusion models is significantly lower compared to Generative Adversarial Networks (GANs)  [14]. Under the diffusion framework, models tend to retain certain images from the training data, potentially generating outputs that closely resemble the original images. To remove explicit artworks from diffusion models, Andikota et al. [7] presents a fine-tuning method for concept removal from diffusion models. Additionally, Zhang et al. [42] presents the “Forget-Me-Not” method, which enables the targeted removal of specific objects and content from diffusion models within a span of 30 seconds, while minimizing the impact on other contents. Somepalli et al. [38] explores whether diffusion models create unique artworks or directly replicate certain contents from the training dataset during image generation. Furthermore, there exist numerous model unlearning methods in the context of image-related tasks, as evidenced by  [3, 11, 16, 15, 22, 26, 8], among others.
@@
Works in Artistic Image Communities. With the emergence of painting capabilities in models like Stable Diffusion, there has been a growing surge in activity and attention within communities focused on image copyright protection. For instance, websites such as stablediffusion.fr/artists and urania.ai/top-sd-artists have gained prominence. In comparison to the efforts of these image communities, our approach involves the collection of real, valuable, and specific images from the art world to be used as training examples for image generation models. This represents a more rigorous form of style imitation. In contrast, the artistic images in the provided links tend to focus on capturing certain aspects of an artist’s style, fitting into a broader category of style imitation. For details, see the supplementary materials.
@@
Policy, Legal, and Social Impact. The increasing global popularity of AIGC highlights the importance of privacy and copyright issues. AI companies, including OpenAI, have taken measures to address concerns related to data security. The US has proposed establishing a new government agency responsible for approving large-scale AI models. Furthermore, the Chinese Cyberspace Administration has published a document emphasizing AIGC security issues [^1] http://www.cac.gov.cn/2023-04/11/c_1682854275475410.htm. Recently enacted legislation, such as the General Data Protection Regulation (GDPR [^2] https://gdpr-info.eu/) in the European Union, the California Consumer Privacy Act (CCPA [^3] https://oag.ca.gov/privacy/ccpa) in California, and the Personal Information Protection and Electronic Documents Act (PIPEDA [^4] https://laws-lois.justice.gc.ca/ENG/ACTS/P-8.6/index.html) in Canada, have legally solidified this right [43, 5].
@@ 3 CPDM Dataset
3.1 Dataset Creation Pipeline
Figure 3: Pipeline for the CPDM Dataset Creation. The clip-interrogator is utilized to convert copyrighted images into corresponding textual information. This text is subsequently refined and transformed into prompts, which are then inputted into a diffusion model to generate the corresponding infringing images.
We propose a pipeline to coordinate CLIP and diffusion models to generate a dataset that contains anchor images, corresponding prompts, and images generated by text-to-image models, reflecting the potential abuses of copyright, as illustrated in Fig. 3. Initially, we collect a set of images that potentially contain copyrighted content, which serves as anchor images. Subsequently, these images are fed into the clip-interrogator [^5] https://github.com/pharmapsychotic/clip-interrogator, allowing us to obtain prompts that correspond to each anchor image. We also provided experimental results using tools other than clip-interrogator in the supplementary materials. Finally, the prompts are used as inputs for the stable diffusion model, resulting in the generation of images by the stable diffusion model. The final outcomes indicate that even such a rudimentary pipeline can effectively generate a substantial volume of works pertinent to infringement issues.